OpenStudy (anonymous):

4 2/3 divided into 275

OpenStudy (anonymous):

\[275\div 4\tfrac{2}{3}\]

OpenStudy (anonymous):

\[275\div\frac{14}{3}\] \[275\times \frac{3}{14}\]